Grade 11Physical Chemistry

Calculate wavelength of a radiation emitted an electron in an hydrogen atom makes a transition from 4th to 2nd energy level

Sukhdev
8 Years agoGrade 11

1 Answer

Arun
8 Years ago

According to Balmer formula

Wave no = RH [1/n12 – 1/n22]

Here n1 = 2 ,       n2 = 4,     RH = 109678

Putting these values in the equation we get

Wave number =109678(1/2- ¼2) = 109678 x 3/16

Also λ = 1/ wave number

Therefore λ = 16 / 109678 x 3 = 486 nm